OpenStudy (anonymous):

solve by elimination 5x=10+2y -4/3x+y=13/3

OpenStudy (skullpatrol):

5x -2y = 10 -4x +3y =13

OpenStudy (anonymous):

5x=10+2y x=(10+2y)/5 Substitute this value of x in 2nd equation.

OpenStudy (anonymous):

(3y-13)/4 =x in place of eq 1 and we get

OpenStudy (anonymous):

-4/3*(10+2y)/5+y=13/3 (-40-8y)/15+15y/15=65/15 -40-8y+15y=65 7y=105 you may calculate y and then x from here. if you have any doubt, please tell.

OpenStudy (anonymous):

yaa thats true